OpenStudy (anonymous):

Evaluate f(x) = x 2 + 1 for f(-1).

ganeshie8 (ganeshie8):

put x = -1

ganeshie8 (ganeshie8):

\(f(x) = x^2 + 1\) to find \(f(-1)\), put \(x =-1\) \(f(-1) = (-1)^2 + 1\) = ?

OpenStudy (anonymous):

you can do it mila

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."